OpenStudy (anonymous):

what the angle in radians between <-5,9,32> and <5,-8,3>

OpenStudy (anonymous):

what is the identity of a dot product, do you know?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Are you in calculus right now?

OpenStudy (anonymous):

Yes calc 3

OpenStudy (anonymous):

What does does \[\huge U \cdot V\] equal?? there is an identity.

OpenStudy (anonymous):

Try this cos theta = v*w/abs(v)*abs(w). So, you can write cos theta= <-5,9,32> x<5,-8,3>/sqrt(-5^2+92+32^2) * sqrt(5^2+(-8)^2+32).

OpenStudy (anonymous):

it's not the absolute value. it's the magnitude, they are not the same.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I don´t know what symbol use here.

OpenStudy (anonymous):

it's alright, just wanted to make sure it was clear so it wasn't confused with doing the wrong thing, ya know?

OpenStudy (anonymous):

Yes, that´s for sure.

OpenStudy (anonymous):

\[\huge \vec U \cdot \vec V = ||\vec U|| * ||\vec V|| \cos(\theta)\]

OpenStudy (anonymous):

ah, don't just give answers.....

OpenStudy (anonymous):

Hey Jessicar make the calculations. Just this.

OpenStudy (anonymous):

But in the formulae above i Think cos theta must written above in the denominator. As we can seen above it seems that we have a multiplication of all entities.